how many minutes would it take a radio wave to travel from the planet venus to earth.? (average distance from venus to earth= 28 million miles) 1 kilometer (km)= 0.6214 mile (mi), 1 km = 1000 m

1
Expert's answer
2022-01-31T10:46:27-0500

Average distance from Venus to earth

"D=28million miles"

"D=28\times10^6\times1.61 km"


"D=45.08\times10^8\times10^3=45.08\times10^9m"

Time

"T=\frac{D}{c}"

"T=\frac{45.08\times10^{9}}{3\times10^{8}}=150.26m\\T=\frac{150.26}{60}=2.54min"